# Computing Distances on Graph Associahedra is Fixed-parameter Tractable

### v1

@@ Abstract
An elimination tree of a connected graph $G$ is a rooted tree on the vertices of $G$ obtained by choosing a root $v$ and recursing on the connected components of $G-v$ to obtain the subtrees of $v$. The graph associahedron of $G$ is a polytope whose vertices correspond to elimination trees of $G$ and whose edges correspond to tree rotations, a natural operation between elimination trees. These objects generalize associahedra, which correspond to the case where $G$ is a path. Ito et al. [ICALP 2023] recently proved that the problem of computing distances on graph associahedra is NP-hard. In this paper we prove that the problem, for a general graph $G$, is fixed-parameter tractable parameterized by the distance $k$. Prior to our work, only the case where $G$ is a path was known to be fixed-parameter tractable. To prove our result, we use a novel approach based on a marking scheme that restricts the search to a set of vertices whose size is bounded by a (large) function of $k$.
@@
graph associahedra, elimination tree, rotation distance, parameterized complexity, fixed-parameter tractable algorithm, combinatorial shortest path, reconfiguration.
@@ 1 Introduction
Given a connected and undirected graph $G$, an elimination tree $T$ of $G$ is any rooted tree that can be defined recursively as follows. If $V(G)=\{v\}$, then $T$ consists of a single root vertex $v$. Otherwise, a vertex $v\in V(G)$ is chosen as the root of $T$, and an elimination tree is created for each connected component of $G-v$. Each root of these elimination trees of $G-v$ is a child of $v$ in $T$. For a disconnected graph $G$, an elimination forest of $G$ is the disjoint union of elimination trees of the connected components of $G$. Equivalently, an elimination forest of a graph $G$ is a rooted forest $F$ (that is, a forest with a root in every connected component) on vertex set $V(G)$ such that for each edge $uv\in E(G)$, vertex $u$ is an ancestor of vertex $v$ in $F$, or vice versa.
@@
Figure 1: A graph $G$ and two of its elimination trees $T$ and $T^{\prime}$, where the second one is obtained from the first one by the rotation of edge $uv$ (in red).
@@
Given a class of combinatorial objects and a “local change” operation between them, the corresponding flip graph has as vertices the combinatorial objects, and its edges connect pairs of objects that differ by the prescribed change operation. In this article, we focus on the case where this class of combinatorial objects is the set of elimination forests of a graph $G$. For these objects, the commonly considered “local change” operation is that of edge rotation defined as follows, where we suppose for simplicity that $G$ is connected. Given an elimination tree $T$ of a graph $G$, the rotation of an edge $uv\in E(T)$, with $u$ being the parent of $v$, creates another elimination of $G$, denoted by ${\sf rot}(T,uv)$, obtained, informally, by just swapping the choice of $u$ and $v$ in the recursive definition of $T$ (that is, in the so-called elimination ordering), and updating the parent of the subtrees rooted at $v$ accordingly; see Figure 2 for an illustration. The formal definition can be found in section 3 (cf. 3.1).
@@
Related work.
@@
Our focus is on the algorithmic problem of determining the distance between two vertices of ${\mathcal{A}}(G)$, or equivalently, determining the rotation distance between two given elimination trees of a graph $G$. There are very few cases where this problem is known to be solvable in polynomial time, namely when $G$ is a complete graph (folklore), a star [9], or a complete split graph [9]. The complexity of the case where $G$ is a path is a notorious long-standing open problem. On the positive side, for $G$ being a path, there exist a polynomial-time $2$-approximation algorithm [15] and several fixed-parameter tractable (FPT) algorithms when the distance is the parameter [14, 25, 26, 28, 30]. It is worth mentioning that there are some hardness results on generalized settings [29, 34, 2] and polynomial-time algorithms for some type of restricted rotations [13].
Cardinal et al. [5] asked whether computing distances on general graph associahedra is NP-hard. Very recently, this question was answered positively by Ito et al. [24].
Our result.
@@
Instance: A graph $G$, two elimination trees $T$ and $T^{\prime}$ of $G$, and a positive integer $k$. Parameter: $k$. Question: Is the rotation distance between $T$ and $T^{\prime}$ at most $k$? Rotation Distance parameterized by
@@
$f(k)=k^{k\cdot 2^{2^{.^{.^{.{2^{{\mathcal{O}}(k^{2})}}}}}}},\text{ where the tower of exponentials has height at most $(3k+1)4k={\mathcal{O}}(k^{2})$.}$
@@
In section 2 we present an overview of the main ideas of the algorithm of Theorem 1.1, which may serve as a road map to read the rest of the article. In section 3 we provide standard preliminaries about graphs and parameterized complexity and fix our notation, in section 4 we formally present our FPT algorithm (split into several subsections), and in section 5 we discuss several directions for further research.
2 Overview of the main ideas of the algorithm
Equipped with the definition of trace, we can define the notion of type, which is somehow involved (cf. 4.9) and whose intuition behind is the following. For our marking algorithm to make sense, we want that if two vertices $v,v^{\prime}$ with the same parent (called $T$-siblings) have the same type (within $Z$), denoted by $\tau(T,Z,v)=\tau(T,Z,v^{\prime})$, and an $\ell$-rotation sequence $\sigma$ from $T$ to $T^{\prime}$ uses some vertex from $T(v)$ but uses no vertex in $T(v^{\prime})$, then there exists another $\ell$-rotation sequence $\sigma^{\prime}$ from $T$ to $T^{\prime}$ that uses vertices in $T(v^{\prime})$ instead of those in $T(v)$. To guarantee this replacement property, we need a stronger condition than just $v$ and $v^{\prime}$ having the same trace. Informally, we need them to have the same “variety of traces among their children within $Z$”. More formally, this leads to a recursive definition where, in the leaves of $Z$ (that are not necessarily leaves of $T$), the type corresponds to the trace, and for non-leaves, the type is defined by the trace and by the number of children of each possible lower type. Note that, a priori, the number of children of a given type may be unbounded, which would rule out the objective of bounding the number of types as a function of $k$. To overcome this obstacle, the crucial and simple observation is that at most $k$ subtrees rooted at a vertex of $T$ contain vertices used by the desired rotation sequence $\sigma$ (cf. 4.14). This implies that if there are at least $k+1$ $T$-siblings of the same type, necessarily the whole subtree of at least one of them, say $u$, will not be used by $\sigma$, implying that $u$ (and its whole subtree) achieves the desired parent in $T^{\prime}$ without being used by $\sigma$, and the same occurs to any other $T$-sibling of the same type. Thus, keeping track of the existence of at least $k+1$ such children (regardless of their actual number) is enough to capture this “static” behavior, and allows us to shrink the possible distinct numbers to keep track of to a function of $k$ (cf. Equation 4, where the “$\min$” is justified by the previous discussion). Finally, for technical reasons we also incorporate into the type of a vertex its desired parent in $T^{\prime}$, in case it defers from its parent in $T^{\prime}$ (cf. function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},\cdot)$). See 4.9 for the formal definition of type and Figure 5 for an example with $k=2$.
@@ 3 Preliminaries
A parameterized problem is a language $L\subseteq\Sigma^{*}\times\mathbb{N}$, for some finite alphabet $\Sigma$. For an instance $(x,k)\in\Sigma^{*}\times\mathbb{N}$, the value $k$ is called the parameter. Such a problem is fixed-parameter tractable (FPT for short) if there is an algorithm that decides membership of an instance $(x,k)$ in time $f(k)\cdot{|x|}^{{\mathcal{O}}(1)}$ for some computable function $f$. Consult [16, 33, 20, 19, 21] for background on parameterized complexity.
@@ 4.1 Good and bad vertices
Our algorithm exploits how a rotation in an elimination tree $T$ may affect the parents and the children of its vertices. Note that a single rotation of an edge $uv\in E(T)$, yielding an elimination tree $T^{\prime}$, may change the parent of arbitrarily many vertices. Indeed, these vertices are the roots of the red subtrees in Figure 2, and the considered vertex $v$ may be adjacent to the root of arbitrarily many subtrees containing at least one vertex adjacent to $u$: for each such root $r$, ${\sf parent}(T,r)=v$ but ${\sf parent}(T^{\prime},r)=u$. As a concrete example, in Figure 1, ${\sf parent}(T,s)=v$ but ${\sf parent}(T^{\prime},s)=u$. On the other hand, item 6 of 3.1 implies that there are at most three vertices whose children set changes from $T$ to $T^{\prime}$, namely $u,v,z$ as depicted in Figure 2. (Note that the sets of children of $u$ and $v$ always change, and that of $z$ changes provided that this vertex exists.) We state this observation formally, since it will be extensively used afterwards.
@@
subsection 4.1 is equivalent to saying that we can safely conclude that any instance $(G,T,T^{\prime},k)$ of Rotation Distance with at least $13k+$ 1 $(T,T^{\prime})$-children-bad vertices is a no-instance. Thus, we can assume henceforth that we are dealing with an instance of Rotation Distance containing at most $3k$ $(T,T^{\prime})$-children-bad vertices.
@@ 4.2 Restricting the rotations to small balls around bad vertices
By 4.3, we focus henceforth on trying to find an $\ell$-rotation sequence from $T$ to $T^{\prime}$, with $\ell\leq k$, consisting only of edges with both endvertices in $B_{\sf cb}$. First, we will consider each of the at most $13k+$ 1 connected components of $T[B_{\sf cb}]$ separately. In fact, we can get a better bound, as if $T[B_{\sf cb}]$ has at least $k+1$ connected components, we can immediately conclude that we are dealing with a no-instance, since at least one rotation is needed in each component. Thus, we may assume that $T[B_{\sf cb}]$ has at most $k$ connected components. On the other hand, since $T[B_{\sf cb}]$ is defined as the union of at most $13k+$ 1 balls of radius $2k$, it follows that every $Z\in{\sf cc}(T[B_{\sf cb}])$ satisfies
@@ 4.3 Description of the marking algorithm
Let $T$ be an elimination tree (of a graph $G$), let $Z$ be a rooted subtree of $T$ corresponding to a connected component of $B_{\sf cb}$, and let $v\in V(Z)$. The trace of $v$ in $Z$, denoted by ${\sf trace}(T,Z,v)$, is a binary vector of dimension ${\sf dist}_{T}(v,{\sf root}(Z))$ defined as follows (note that if $v={\sf root}(Z)$, then its trace is empty). For $i\in[{\sf dist}_{T}(v,{\sf root}(Z))]$, let $u_{i}\in{\sf ancestors}(T,v)$ be the ancestor of $v$ in $T$ such that ${\sf dist}_{T}(v,u_{i})=i$. Then the $i$-th coordinate of ${\sf trace}(T,Z,v)$ is $1$ if $wu_{i}\in E(G)$ for some vertex $w\in V(T(v))$, and 00 otherwise.
@@
For a vertex $v\in V(T)$, let ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$ be equal to $\emptyset$ if ${\sf parent}(T,v)={\sf parent}(T^{\prime},v)$, and to ${\sf parent}(T^{\prime},v)$ otherwise. Note that, by subsection 4.1, the function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$ can take up to $13k+$ 1 distinct values when ranging over all $v\in V(T)$.
@@
Figure 5: A component $Z$ of $T[B_{\sf cb}]$ and the types of its vertices, for an instance with $k=2$. For the sake of simplicity, different types are depicted with different numbers. Assume that the leaves have only two possible types, namely $1$ and $2$, and that all non-leaf vertices at the same distance from the root have the same trace and the same function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},\cdot)$. Note that the red vertices have the same type (namely, $4$) because they both have one child of type $1$ and at least $k+1=3$ children of type $2$. Note also that the blue vertices have the same type (namely, $6$) because they both have one child of type $3$ and one child of type $4$.
@@
$g(k)=k^{2^{2^{.^{.^{.{2^{{\mathcal{O}}(k^{2})}}}}}}},\text{ where the tower ofexponentials has height ${\sf diam}(Z)={\mathcal{O}}(k^{2})$.}$ (5)
@@
By the definition of type, $\tau_{d}\leq 2^{d}\cdot(3k+1)$ since, on the one hand, all vertices at distance $d$ from ${\sf root}(Z)$ are leaves of $Z$, and the number of possible traces among leaves is at most $2^{d}$, and on the other hand the term $13k+$ 1 corresponds to the possible distinct values of the function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$. For every $i\in[d-1]$, Equation 4 implies that
@@
where the term $13k+$ 1 again comes from the possible distinct values of the function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$, the term $2^{i}$ comes from the possible different traces within distance $i$ from ${\sf root}(Z)$, and the term $k+2$ follows from the fact that, for every $v\in V(Z)$, the function $f_{v}$ can take up to $k+1$ values for each type $\tau$ of a children of $v$, together with the possibility that a type is not present among the children of $v$. Note that a vertex $v\in V(T)$ with ${\sf dist}(v,{\sf root}(Z))=i$ may have children being roots of any possible subtree with diameter at most $d-i$, justifying the sum in Equation 7. Clearly, the upper bound of Equation 7 is maximized for $i=1$, that is, for the children of ${\sf root}(Z)$, yielding the bound claimed in Equation 5.
@@
  • If |$|\{u\in{\sf children}(Z,v)\mid\tau(Z,u)=\tau\}|\leq k+1$, add the whole set $\{u\in{\sf children}(Z,v)\mid\tau(Z,u)=\tau\}$ to $M_{Z}^{\sf pre}$.
@@
Let us analyze the size of each set $M_{Z}$ of $T[M]$ separately, as their number is at most $13k+$ 1, and this factor gets subsumed by the asymptotic growth of $h(k)$. The diameter of $T[M_{Z}]$ is at most $(3k+1)4k={\mathcal{O}}(k^{2})$ by Equation 3, so in order to bound the size of $M_{Z}$, it just remains to bound the degree of $T[M_{Z}]$. By construction of $M_{Z}$, every vertex $v\in M_{Z}$ has at most ${\sf children}(Z,v)\cap M_{Z}^{\sf pre}+3k$ children, where the term $3k$ corresponds to the maximum number of $(T,T^{\prime})$-children-bad vertices (cf. subsection 4.1) that, together with their ancestors, have been marked within $Z$. Since the set $M_{Z}^{\sf pre}$ is defined by pre-marking, for each vertex, at most $k+1$ children of each type, the set ${\sf children}(Z,v)\cap M_{Z}^{\sf pre}$ has size at most $g(k)\cdot(k+1)$, where $g(k)$ is the function given by 4.10. Thus, $|M|\leq h(k)$, where $h(k)$ has the same asymptotic growth as $g(k)$.

### v2

# Computing distances is FPT on graph associahedra and W[2]-hard on hypergraphic polytopes

@@ Abstract
An elimination tree of a connected graph $G$ is a rooted tree on the vertices of $G$ obtained by choosing a root $v$ and recursing on the connected components of GvG-v to obtain the subtrees of $v$. The graph associahedron of $G$ is a polytope whose vertices correspond to elimination trees of $G$ and whose edges correspond to tree rotations, a natural operation between elimination trees. These objects generalize associahedra, which correspond to the case where $G$ is a path. Ito et al. [SIAM J. Discrete Math. 2026] recently proved that the problem of computing distances on graph associahedra is NP-hard. In this paper we prove that the problem, for a general graph $G$, is fixed-parameter tractable (FPT) parameterized by the distance $k$. Prior to our work, only the case where $G$ is a path was known to be FPT. To prove our result, we use a novel approach based on a marking scheme that restricts the search to a set of vertices whose size is bounded by a (large) function of $k$.
On the negative side, we show that it is unlikely that FPT algorithms exist on a natural generalization of graph associahedra, namely hypergraphic polytopes, by proving that computing distances on them is ${\sf W}[2]$-hard parameterized by the distance. We also prove that, on hypergraphic polytopes, the distance cannot be approximated in polynomial time within a factor $c\cdot\log(|V|+|\mathcal{E}|)$ for some constant $c>0$ unless ${\sf P}={\sf NP}$, where $H=(V,\mathcal{E})$ is the input hypergraph. This result strengthens the hardness result of Cardinal and Steiner [Combin. Theory 2025], who proved that the problem cannot be approximated within a factor $(1+\varepsilon)$ for some absolute constant $\varepsilon>0$ unless ${\sf P}={\sf NP}$. Finally, we rule out the existence of polynomial kernels parameterized by the number of vertices of the input hypergraph, a parameter for which the problem is easily seen to be FPT.
@@
graph associahedra, elimination tree, rotation distance, parameterized complexity, fixed-parameter tractable algorithm, combinatorial shortest path, reconfiguration, hypergraphic polytopes, orientation of hypergraphs, hardness of approximation, polynomial kernels.
@@ 1 Introduction
Given a connected and undirected graph $G$, an elimination tree $T$ of $G$ is any rooted tree that can be defined recursively as follows. If $V(G)=\{v\}$, then $T$ consists of a single root vertex $v$. Otherwise, a vertex $v\in V(G)$ is chosen as the root of $T$, and an elimination tree is created for each connected component of GvG-v. Each root of these elimination trees of GvG-v is a child of $v$ in $T$. For a disconnected graph $G$, an elimination forest of $G$ is the disjoint union of elimination trees of the connected components of $G$. Equivalently, an elimination forest of a graph $G$ is a rooted forest $F$ (that is, a forest with a root in every connected component) on vertex set $V(G)$ such that for each edge $uv\in E(G)$, vertex $u$ is an ancestor of vertex $v$ in $F$, or vice versa.
@@
Figure 1: A graph $G$ and two of its elimination trees $T$ and $T^{\prime}$, where the second one is obtained from the first one by the rotation of edge uvuv (in red).
@@
Given a class of combinatorial objects and a “local change” operation between them, the corresponding flip graph has as vertices the combinatorial objects, and its edges connect pairs of objects that differ by the prescribed change operation. In this article, we focus on the case where this class of combinatorial objects is the set of elimination forests of a graph $G$. For these objects, the commonly considered “local change” operation is that of edge rotation defined as follows, where we suppose for simplicity that $G$ is connected. Given an elimination tree $T$ of a graph $G$, the rotation of an edge $uv\in E(T)$, with $u$ being the parent of $v$, creates another elimination of $G$, denoted by ${\sf rot}(T,uv)$, obtained, informally, by just swapping the choice of $u$ and $v$ in the recursive definition of $T$ (that is, in the so-called elimination ordering), and updating the parent of the subtrees rooted at $v$ accordingly; see Figure 2 for an illustration. The formal definition can be found in section 2 (cf. 2.1).
@@
Related work on graph associahedra.
@@
Our focus is on the algorithmic problem of determining the distance between two vertices of ${\mathcal{A}}(G)$, or equivalently, determining the rotation distance between two given elimination trees of a graph $G$. There are very few cases where this problem is known to be solvable in polynomial time, namely when $G$ is a complete graph (folklore), a star [12], or a complete split graph [12]. The complexity of the case where $G$ is a path has been a notorious long-standing open problem for several decades, and it has been very recently settled by Dorfer [23], who has proved it to be NP-complete. On the positive side, for $G$ being a path, there exist a polynomial-time $2$-approximation algorithm [18] and several fixed-parameter tractable (FPT) algorithms when the distance is the parameter [17, 32, 33, 35, 37]. It is worth mentioning that there are some hardness results on generalized settings [36, 41, 3] and polynomial-time algorithms for some type of restricted rotations [16].
Cardinal et al. [8] asked whether computing distances on general graph associahedra is NP-hard (before the hardness result of Dorfer [23] was known). This question was answered positively by Ito et al. [31].
Our results.
@@
Instance: A graph $G$, two elimination trees $T$ and $T^{\prime}$ of $G$, and a positive integer $k$.
Parameter: $k$.
Question: Is the rotation distance between $T$ and $T^{\prime}$ at most $k$?
Rotation Distance
@@
[…]
@@
It is natural to ask whether the FPT algorithm of Theorem 1.1 can be generalized beyond graph associahedra. In the second part of the article we prove that, for a natural generalization of graph associahedra called hypergraphic polytopes (defined in subsection 5.1, see [13, 4, 29, 52, 2, 44, 43, 48] for more background), this is probably not the case, since we prove (cf. Theorem 5.2) that computing distances on hypergraphic polytopes is ${\sf W}[2]$-hard parameterized by the distance. We prove this by a parameterized reduction from Dominating Set parameterized by the size of the solution.
By appropriately modifying the ${\sf W}[2]$-hardness reduction, we also transfer the hardness of approximation from Dominating Set to computing distances on hypergraphic polytopes. Namely, as it is the case of Dominating Set [47], we prove (cf. Theorem 5.8) that there exists some constant $c>0$ such that computing distances on hypergraphic polytopes does not admit a polynomial-time $c\cdot\log(|V|+|{\mathcal{E}}|)$-approximation algorithm unless ${\sf P}={\sf NP}$, where $H=(V,{\mathcal{E}})$ is the input hypergraph. This result strengthens the recent hardness result of Cardinal and Steiner [13], who proved that the problem cannot be approximated within a factor $(1+\varepsilon)$ for some absolute constant $\varepsilon>0$ unless ${\sf P}={\sf NP}$.
Finally, another appropriate modification of the ${\sf W}[2]$-hardness reduction, namely by reducing from the Red-Blue Dominating Set problem instead, allows us (cf. Theorem 5.11) to rule out the existence of polynomial kernels for computing distances on hypergraphic polytopes parameterized by the number of vertices of the input hypergraph, unless ${\sf NP}\subseteq{\sf coNP}/{\sf poly}$. As we discuss in subsection 5.4, the problem is easily seen to be FPT with this parameterization.
@@
In section 2 we provide standard preliminaries about graphs and parameterized complexity and fix our notation. In section 3 we present an overview of the main ideas of the algorithm of Theorem 1.1, which may serve as a road map to read the formal description of the FPT algorithm presented in section 4, split into several subsections. We prove the hardness results in section 5, and in section 6 we discuss several directions for further research, both on graph associahedra and hypergraphic polytopes.
@@ 2 Preliminaries
A parameterized problem is a language $L\subseteq\Sigma^{*}\times\mathbb{N}$, for some finite alphabet $\Sigma$. For an instance $(x,k)\in\Sigma^{*}\times\mathbb{N}$, the value $k$ is called the parameter. Such a problem is fixed-parameter tractable (FPT for short) if there is an algorithm that decides membership of an instance $(x,k)$ in time $f(k)\cdot{|x|}^{{\mathcal{O}}(1)}$ for some computable function $f$.
For a computable function $g\colon\mathbb{N}\to\mathbb{N}$, a kernelization algorithm (or simply a kernel) for a parameterized problem $L$ of size $g$ is an algorithm $A$ that given any instance $(x,k)$ of $L$, runs in polynomial time and returns an instance $(x^{\prime},k^{\prime})$ such that $(x,k)\in L\Leftrightarrow(x^{\prime},k^{\prime})\in L$ with $|x^{\prime}|$, $k^{\prime}\leq g(k)$. The function $g(k)$ is called the size of the kernel, and a kernel is polynomial (resp. linear, quadratic) if $g(k)$ is a polynomial (resp. linear, quadratic) function.
To transfer the non-existence of polynomial kernels (under reasonable complexity assumptions), we use the notion of polynomial parameter transformation (PPT for short), introduced by Bodlaender, Thomassé, and Yeo [6]. A polynomial parameter transformation from a parameterized problem $P$ to a parameterized problem $Q$ is an algorithm that, given an instance $(x,k)$ of $P$, computes in polynomial time an equivalent instance $(x^{\prime},k^{\prime})$ of $Q$ such that $k^{\prime}$ is bounded by a polynomial depending only on $k$. It follows easily from the definition that if $P$ does not admit a polynomial kernel under some complexity assumption, then the same holds for $Q$.
Within parameterized problems, the classes ${\sf W}[i]$ with $i\geq 1$ may be seen as the parameterized equivalent to the class NP of classical decision problems. Without entering into details (see [24, 19] for the formal definitions), a parameterized problem being W[i]-hard for some $i\geq 1$ can be seen as a strong evidence that this problem is not FPT. The canonical example of ${\sf W}[1]$-hard (resp. ${\sf W}[2]$-hard) problem is Clique (resp. Dominating Set), both parameterized by the size of the solution
Consult [19, 40, 26, 24, 27] for background on parameterized complexity.
3 Overview of the main ideas of the algorithm
Equipped with the definition of trace, we can define the notion of type, which is somehow involved (cf. 4.9) and whose intuition behind is the following. For our marking algorithm to make sense, we want that if two vertices $v,v^{\prime}$ with the same parent (called $T$-siblings) have the same type (within $Z$), denoted by $\tau(T,Z,v)=\tau(T,Z,v^{\prime})$, and an $\ell$-rotation sequence $\sigma$ from $T$ to $T^{\prime}$ uses some vertex from $T(v)$ but uses no vertex in $T(v^{\prime})$, then there exists another $\ell$-rotation sequence $\sigma^{\prime}$ from $T$ to $T^{\prime}$ that uses vertices in $T(v^{\prime})$ instead of those in $T(v)$. To guarantee this replacement property, we need a stronger condition than just $v$ and $v^{\prime}$ having the same trace. Informally, we need them to have the same “variety of traces among their children within ZZ”. More formally, this leads to a recursive definition where, in the leaves of $Z$ (that are not necessarily leaves of $T$), the type corresponds to the trace, and for non-leaves, the type is defined by the trace and by the number of children of each possible lower type. Note that, a priori, the number of children of a given type may be unbounded, which would rule out the objective of bounding the number of types as a function of $k$. To overcome this obstacle, the crucial and simple observation is that at most $k$ subtrees rooted at a vertex of $T$ contain vertices used by the desired rotation sequence $\sigma$ (cf. 4.14). This implies that if there are at least $k+1$ $T$-siblings of the same type, necessarily the whole subtree of at least one of them, say $u$, will not be used by $\sigma$, implying that $u$ (and its whole subtree) achieves the desired parent in $T^{\prime}$ without being used by $\sigma$, and the same occurs to any other $T$-sibling of the same type. Thus, keeping track of the existence of at least $k+1$ such children (regardless of their actual number) is enough to capture this “static” behavior, and allows us to shrink the possible distinct numbers to keep track of to a function of $k$ (cf. Equation 4, where the $\min$” is justified by the previous discussion). Finally, for technical reasons we also incorporate into the type of a vertex its desired parent in $T^{\prime}$, in case it defers from its parent in $T^{\prime}$ (cf. function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},\cdot)$). See 4.9 for the formal definition of type and Figure 5 for an example with $k=2$.
@@ 4.1 Good and bad vertices
Our algorithm exploits how a rotation in an elimination tree $T$ may affect the parents and the children of its vertices. Note that a single rotation of an edge $uv\in E(T)$, yielding an elimination tree $T^{\prime}$, may change the parent of arbitrarily many vertices. Indeed, these vertices are the roots of the red subtrees in Figure 2, and the considered vertex $v$ may be adjacent to the root of arbitrarily many subtrees containing at least one vertex adjacent to $u$: for each such root $r$, ${\sf parent}(T,r)=v$ but ${\sf parent}(T^{\prime},r)=u$. As a concrete example, in Figure 1, ${\sf parent}(T,s)=v$ but ${\sf parent}(T^{\prime},s)=u$. On the other hand, item 6 of 2.1 implies that there are at most three vertices whose children set changes from $T$ to $T^{\prime}$, namely $u,v,z$ as depicted in Figure 2. (Note that the sets of children of $u$ and $v$ always change, and that of $z$ changes provided that this vertex exists.) We state this observation formally, since it will be extensively used afterwards.
@@
subsection 4.1 is equivalent to saying that we can safely conclude that any instance $(G,T,T^{\prime},k)$ of Rotation Distance with at least $3k+1$ $(T,T^{\prime})$-children-bad vertices is a no-instance. Thus, we can assume henceforth that we are dealing with an instance of Rotation Distance containing at most $3k$ $(T,T^{\prime})$-children-bad vertices.
@@ 4.2 Restricting the rotations to small balls around bad vertices
By 4.3, we focus henceforth on trying to find an $\ell$-rotation sequence from $T$ to $T^{\prime}$, with $\ell\leq k$, consisting only of edges with both endvertices in $B_{\sf cb}$. First, we will consider each of the at most $3k+1$ connected components of $T[B_{\sf cb}]$ separately. In fact, we can get a better bound, as if $T[B_{\sf cb}]$ has at least $k+1$ connected components, we can immediately conclude that we are dealing with a no-instance, since at least one rotation is needed in each component. Thus, we may assume that $T[B_{\sf cb}]$ has at most $k$ connected components. On the other hand, since $T[B_{\sf cb}]$ is defined as the union of at most $3k+1$ balls of radius $2k$, it follows that every $Z\in{\sf cc}(T[B_{\sf cb}])$ satisfies
@@ 4.3 Description of the marking algorithm
Let $T$ be an elimination tree (of a graph $G$), let $Z$ be a rooted subtree of $T$ corresponding to a connected component of $B_{\sf cb}$, and let $v\in V(Z)$. The trace of $v$ in $Z$, denoted by ${\sf trace}(T,Z,v)$, is a binary vector of dimension ${\sf dist}_{T}(v,{\sf root}(Z))$ defined as follows (note that if $v={\sf root}(Z)$, then its trace is empty). For $i\in[{\sf dist}_{T}(v,{\sf root}(Z))]$, let $u_{i}\in{\sf ancestors}(T,v)$ be the ancestor of $v$ in $T$ such that ${\sf dist}_{T}(v,u_{i})=i$. Then the $i$-th coordinate of ${\sf trace}(T,Z,v)$ is $1$ if $wu_{i}\in E(G)$ for some vertex $w\in V(T(v))$, and 0 otherwise.
@@
For a vertex $v\in V(T)$, let ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$ be equal to $\emptyset$ if ${\sf parent}(T,v)={\sf parent}(T^{\prime},v)$, and to ${\sf parent}(T^{\prime},v)$ otherwise. Note that, by subsection 4.1, the function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$ can take up to $3k+1$ distinct values when ranging over all $v\in V(T)$.
@@
Figure 5: A component $Z$ of $T[B_{\sf cb}]$ and the types of its vertices, for an instance with $k=2$. For the sake of simplicity, different types are depicted with different numbers. Assume that the leaves have only two possible types, namely 11 and $2$, and that all non-leaf vertices at the same distance from the root have the same trace and the same function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},\cdot)$. Note that the red vertices have the same type (namely, 44) because they both have one child of type 11 and at least $k+1=3$ children of type 22. Note also that the blue vertices have the same type (namely, 66) because they both have one child of type $3$ and one child of type 44.
@@
[…]
@@
By the definition of type, $\tau_{d}\leq 2^{d}\cdot(3k+1)$ since, on the one hand, all vertices at distance $d$ from ${\sf root}(Z)$ are leaves of $Z$, and the number of possible traces among leaves is at most $2^{d}$, and on the other hand the term $3k+1$ corresponds to the possible distinct values of the function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$. For every $i\in[d-1]$, Equation 4 implies that
@@
where the term $3k+1$ again comes from the possible distinct values of the function ${\sf{\sf want{\text{-}}parent}}(T,T^{\prime},v)$, the term $2^{i}$ comes from the possible different traces within distance $i$ from ${\sf root}(Z)$, and the term $k+2$ follows from the fact that, for every $v\in V(Z)$, the function $f_{v}$ can take up to $k+1$ values for each type $\tau$ of a children of $v$, together with the possibility that a type is not present among the children of $v$. Note that a vertex $v\in V(T)$ with ${\sf dist}(v,{\sf root}(Z))=i$ may have children being roots of any possible subtree with diameter at most did-i, justifying the sum in Equation 7. Clearly, the upper bound of Equation 7 is maximized for $i=1$, that is, for the children of ${\sf root}(Z)$, yielding the bound claimed in Equation 5.
@@
  • If $|\{u\in{\sf children}(Z,v)\mid\tau(Z,u)=\tau\}|\leq k+1$, add the whole set $\{u\in{\sf children}(Z,v)\mid\tau(Z,u)=\tau\}$ to $M_{Z}^{\sf pre}$.
@@
Let us analyze the size of each set $M_{Z}$ of $T[M]$ separately, as their number is at most $3k+1$, and this factor gets subsumed by the asymptotic growth of $h(k)$. The diameter of $T[M_{Z}]$ is at most $(3k+1)4k={\mathcal{O}}(k^{2})$ by Equation 3, so in order to bound the size of $M_{Z}$, it just remains to bound the degree of $T[M_{Z}]$. By construction of $M_{Z}$, every vertex $v\in M_{Z}$ has at most ${\sf children}(Z,v)\cap M_{Z}^{\sf pre}+3k$ children, where the term $3k$ corresponds to the maximum number of $(T,T^{\prime})$-children-bad vertices (cf. subsection 4.1) that, together with their ancestors, have been marked within $Z$. Since the set $M_{Z}^{\sf pre}$ is defined by pre-marking, for each vertex, at most $k+1$ children of each type, the set ${\sf children}(Z,v)\cap M_{Z}^{\sf pre}$ has size at most $g(k)\cdot(k+1)$, where $g(k)$ is the function given by 4.10. Thus, $|M|\leq h(k)$, where $h(k)$ has the same asymptotic growth as $g(k)$.